OpenStudy (anonymous):

7x + 2(x-4) = 3x-6 + 2x, I got 10/9 is that right?

OpenStudy (anonymous):

No 7x + 2(x-4) = 3x-6 + 2x 7x+2x-8=5x-6 9x-5x-8=-6 4x=-8+8 4x=2 x=1/2

OpenStudy (anonymous):

Thanks